OpenStudy (anonymous):

Which of the following systems of equations has infinitely many solutions? 3x + 3y = 11; 6y = 6x − 22 8x − 7 = 6y; 12y − 14 = 4x 7y = x + 3; -3x + 21y = 9 5x − 7y = 4; 14y = 10x + 8

OpenStudy (anonymous):

@dpasingh can you help me with this please

OpenStudy (anonymous):

can you help me with the next one

OpenStudy (anonymous):

5x − 7y = 4; 14y = 10x + 8

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve using substitution: 2y − x + 2 = 0 -3x + 8y = -5 (-10, 4) (3, 1/2) (2, 0) (4, 1)

OpenStudy (anonymous):

Ok wait!!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2y − x + 2 = 0 i.e. 2y+2=x----1 -3x + 8y = -5-------2 substituting x= 2y+2 in eq 2 we find -3( 2y+2) +8y=-5 -6y-6+8y=-5 2y-6=-5 2y=6-5 2y=1 y=1/2 substituting y=1/2 in eq --1 we find x= 2y+2=2(1/2)+2=1+2=3 Thus (x,y)=(3, 1/2) is the correct ordered pair